yielding 4 percent per year, and there are 365 days per year. MINICASE Cash Management at Webb Corporation Webb Corporation w1.   What is Webb Corporation’s total net cash flow available from the current lockbox system to meet payroll?
2.   Under the terms outlined by Third National Bank, should the company proceed with the concentration banking system?
3.   What cost of ACH transfers would make the company indifferent between the two systems?
Answer #1

1. Net cash flow from Current lock-box system:

Webb corporation has 4 collection centres and lockbox at each collection centre handles $ 193,000 in payment each day. Until two weeks (i..e) 15 days, money remains in short term marketable security which earns .012% of income and .20% as wire transfer charges.

Cash inflow = 193000*4centres*15days = $11,580,000

Cash infow = 193000*4centres*15days*0.012% = $ 1,390

Cash outflow = 193000*4 centres*15days*0.20% = $ 23,160

Nett Cash flow = $ 11,558,230 is available to meet with payroll

2. Proposal from Third National bank:

Interest can be earned one day short (i.e.) on the day fund received, interest will not be earned and ACH charges will be $ 150 per transfer. So, Cash outflow would be $ 193000 from 4 centres charged at $ 150 per transfer, which means $ 600 will be the charges per day as against $ 1544 in current lock box system (i.e.) $ 23160 for 15 days, so per day = $ 1544.

Interest income short by 93 (i.e.) $ 1390 for 15 days becomes $ 93 per day whereas Transfer charges decrease by $944. Thus net cash flow increase by $851.

Also, another major advantage is two week wait time in current lock box system, whereas fund available to use from 2nd day of collection in case of concentrated banking system.

So, it is good idea to proceed with concentration banking system proposal from Third National Bank.

3. Difference in cost of ACH Transfer:

Cash outflow in current lock box system is .20% of amount transferred and not fixed amount. Considering business growth, if collection increases in future, the charges will also increase, which results in increased cost for business. As against the same, Third National bank proposal has fixed charges for transfer. So irrespective of the transfer amount, this charges are fixed and easy to project future expense of the business.